Solve the system:
10x+y=-21
Y=5x-6

Found 2 solutions by ewatrrr, Alan3354:
Answer by ewatrrr(24785)

 
Hi
Solve the system: Will use substtution
10x+y=-21
Y=5x-6
10x + (5x-6) = -21
15x = -15
x = -1 and y = -11 (5·-1 -6)
CHECKING our Answer***
-10 + (-11) = -21

Answer by Alan3354(69443)

Solve the system:
10x+y=-21
y=5x-6
----------
10x+y=-21
Sub for y
10x + (5x-6) = -21
15x = -15
x = -1
----------
y = -11